# Conditional Autoencoder for Generating BNS Waveforms with Tidal and Precession Effects

### v1

@@ III.2 Architecture and Hyperparameters of the Model
Our study employs the cAEs to model the amplitude and phase of GW waveforms from BNS mergers. The model consists of two independent cAE, each responsible for learning a low-dimensional representation of either the amplitude $y_{A}$ or phase $y_{\Phi}$ and reconstructing waveforms conditioned on $\Theta$. Each cAE comprises two encoders (Encoder 1 and Encoder 2) and a decoder. Encoder 1 processes the amplitude or phase data, while Encoder 2 encodes the $\Theta$, and the two latent representations are combined in the latent space before being mapped back to a complete waveform by the decoder. The overall model architecture is shown in FIG. 3, where Training and Test sections correspond to the training and inference workflows, while the remaining sections detail the structural components.
Figure 3: Training and testing framework of the our model, along with detailed structures of individual components, the structure of the Transformer and ResNet block are shown in Appendix A (FIG 10).
As shown in FIG. 3, our framework consists of two main branches: one for waveform encoding and reconstruction, and one for encoding $\Theta$. The goal is to align latent representations from both branches while ensuring accurate waveform reconstruction.
@@
Latent variables and loss functions: During training, both the waveform and the $\Theta$ are encoded into their respective latent representations, denoted as $z_{A/\Phi}$ and $z_{\Theta}$. To ensure that the decoder can accurately reconstruct the input waveform and that both latent spaces are aligned, we define two loss components: First, the reconstruction loss is computed as the MAE between the input waveform $x$ and the reconstructed waveform $\hat{h}$:
@@
The total loss is then defined as a weighted sum of the two as Eq. (11):
@@ IV Training and Verification
Our model is trained using the Adam optimizer with an initial learning rate of $10^{-4}$. To enhance training stability and convergence speed, the following learning rate scheduling mechanism is implemented: if the validation loss (val_loss) does not improve for 7 consecutive epochs, the learning rate is reduced by a factor of 0.7, with a minimum threshold of $10^{-8}$. Additionally, an early stopping strategy is employed to prevent overfitting—if the validation loss shows no significant improvement within 15 epochs, training is automatically halted, and the model parameters are reverted to those corresponding to the best validation performance. The training is conducted on a single NVIDIA RTX 3090 GPU, with Amplitude ($A$) and phase ($\Phi$) models trained separately. The total training process took approximately one week on a single RTX 3090. The dataset consists of $3\times 10^{5}$ samples, with 90% used for training and 10% for validation. The batch size is set to 10, and the maximum training epochs are limited to 500. The training of both the amplitude and phase models took approximately a total of seven days. FIG. 4 illustrates the training loss evolution for both the amplitude and phase models.
@@
To evaluate the accuracy of cAE-generated waveforms, we compute the mismatch between the model-generated waveforms and the IMRPhenomXP_NRTidalv2 target waveforms. The evaluation is based on two key metrics: Overlap [54, 55, 56] and Mismatch, which quantify the waveform reconstruction quality. The analysis is conducted in the frequency domain after applying a Fourier transform to the time-domain waveforms, allowing for a more effective comparison of waveform similarities. In GW data analysis, the inner product of two waveforms is typically defined as a noise-weighted integral over the frequency domain, incorporating the power spectral density $S_{n}(f)$:
[…]
where $\tilde{h}_{1}(f)$ and $\tilde{h}_{2}(f)$ are the Fourier transforms of $h_{1}(t)$ and $h_{2}(t)$, respectively, and $S_{n}(f)$ represents the power spectral density (PSD) of the detector noise. This weighted inner product provides a measure of how well two waveforms match in the presence of detector noise. Since our task does not require a realistic noise model, we set $S_{n}(f)=1$.
@@ V Results and Analysis
In this section, we present the accuracy and efficiency of the model on the test set. Furthermore, we analyze various factors that influence the precision of the generated waveforms.
@@ V.1 Mismatch Evaluation
To assess the accuracy of cAE-generated waveforms, we regenerated $3\times 10^{5}$ test samples based on the $\Theta$ ranges listed in TABLE 1 and computed their mismatch values .
Figure 5: The mismatch distribution of the test sample.
FIG. 5 presents the mismatch distribution for $h_{+}$ and $h_{\times}$, showing that their averaged mismatch values are $3.9717\times 10^{-3}$ and $3.9770\times 10^{-3}$, respectively, corresponding to an average waveform accuracy of 99.6%, considering both precession and tidal effects in BNS. Some more detailed charts illustrating the reconstruction results for samples with different orbital cycle numbers are provided in FIG. 11-FIG. 14 in the Appendix A.
To further analyze the factors influencing mismatch, we computed the orbital cycle number for each sample. The time-domain signal $s(t)$ undergoes a Hilbert transform to obtain its analytic representation[57]:
@@
$\text{NumCycles}=\frac{\Delta\phi}{2\pi}.$ (22)
FIG. 6(a) shows the test set’s cycle number distribution with a averaged cycle number of 265 and 68% of the samples falling within the range of 231 to 307 cycles.
Figure 6: (a) The test set’s cycle number distribution. (b) Relationship between the number of cycles and mismatch (log10) in the test set. The color bar indicates the number of samples in each hexagonal bin.
As shown in FIG. 6(b), most samples cluster around 250 cycles, where the mismatch remains relatively low. However, when the cycle number exceeds 300, the mismatch begins to increase. This trend, consistent with FIG. 6(b), indicates that an increasing cycle number significantly impacts the mismatch, particularly when exceeding 300, where the model’s reconstruction accuracy deteriorates.
Figure 7: (a) Relationship between binary mass and cycle number in the test set. (b) Relationship between binary mass, cycle number, and mismatch. (c) Relationship between tidal deformability and mismatch in the test set. (d) Relationship between spin parameters and mismatch.
As shown in FIG. 7(a), lower-mass binaries exhibit more waveform cycles, whereas increasing mass results in fewer cycles, so that mismatch values are relatively high in both low-mass (high-cycle) and high-mass (low-cycle) regions (as shown in FIG. 7(b)). This trend is primarily due to two factors: (1) as the number of waveform cycles increases, the time-domain complexity of the signal rises, making it more challenging for the network to learn; (2) high-cycle samples are underrepresented in the training set, reducing the model’s generalization ability in this region. When the cycle number falls between 200 and 250, the mismatch is approximately $3.1\times 10^{-3}$, increasing slightly to $3.2\times 10^{-3}$ in the 250–300 range. However, a significant rise to $7.2\times 10^{-3}$ is observed for 300–350 cycles, and once the cycle number exceeds 350, the mismatch escalates sharply, reaching $4.6\times 10^{-2}$ in the 350–400 range.
In order to investigate the impact of tidal deformation and spin parameters on mismatch, we plotted the mismatch distributions under different tidal deformability and total spin conditions in FIG. 7. FIG. 7(c) shows that the mismatch distribution under different tidal deformabilities is random, which suggests that, under the conditions of this study, tidal deformability has a negligible impact on waveform generation accuracy. Similarly, FIG. 7(d) indicates that spin parameters also exhibit a weak correlation with mismatch, implying that precession effects do not significantly degrade the model’s waveform accuracy. This observation is further supported by FIG. 15 in the Appendix A, where no strong mismatch trend is observed in the joint distributions of spin and tidal deformability parameters.
Figure 8: Relationship between average mismatch and total mass in the test set.
Finally, we analyze the relationship between total mass and mismatch, as shown in FIG. 8. The results indicate that samples with total masses between $3.0M_{\odot}$ and $5.0M_{\odot}$ exhibit the lowest average mismatch. That is consistent with FIG. 7(b), further confirming that our model achieves the highest accuracy when generating waveforms with total binary masses between $3.0M_{\odot}$ and $5.0M_{\odot}$ and cycles between 200 and 250. Additionally, the joint distribution of binary mass in FIG. 15 of the Appendix A reveals a clear trend of increasing mismatch in both high-mass and low-mass regions.
V.2 Waveform Generation Efficiency
To evaluate the efficiency of waveform generation, we compared the cAE model with several traditional waveform approximation methods by measuring the total computation time for different batch size. Specifically, we selected four widely used models: SpinTaylorT1, IMRPhenomPv2_NRTidal, IMRPhenomPv2_NRTidalv2, and IMRPhenomXP_NRTidalv2. All experiments were conducted using an NVIDIA RTX 3090 GPU and an Intel(R) Xeon(R) Silver 4214R CPU, with batch size ranging from 1 to 10,000. The results indicate that the cAE significantly outperforms traditional methods in computational efficiency, particularly for large-scale batch waveform generation.
For generating a single waveform, the cAE requires $0.12$ s, which is shorter than SpinTaylorT1 ($0.29$ s), IMRPhenomPv2_NRTidalv2 [38] ($0.39$ s), IMRPhenomXP_NRTidalv2 ($0.38$ s), and IMRPhenomPv2_NRTidal [58] ($0.62$ s). For the large-scale generation of 10,000 waveforms, the cAE completes the task in only $7.48$ s, while SpinTaylorT1, IMRPhenomXP_NRTidalv2, IMRPhenomPv2_NRTidalv2, and IMRPhenomPv2_NRTidal require $486.86$, $454.51$, $462.87$, and $497.75$ s, respectively. This indicates that our model can efficiently accelerate waveform generation when generating large batch size. A detailed comparison of computation times is presented in FIG. 9 and TABLE 2. Further analysis of total computation time across varying batch size reveals that the cAE exhibits near-sublinear scaling behavior. As the batch size increases from $1$ to $10$, $100$, $500$, $1{,}000$, and $10{,}000$, the total generation time modestly rises from $0.12$ s to $0.12$, $0.20$, $0.50$, $0.86$, and $7.48$ s, respectively. This scaling highlights the excellent efficiency and scalability of the cAE model for batch processing.
In contrast, traditional waveform approximation methods exhibit significantly steeper increases in total computation time as batch size grow. Specifically, SpinTaylorT1 computation time rises from $0.29$ s (batch size = 1) to $486.86$ s (batch size = 10,000). IMRPhenomPv2_NRTidal grows from $0.62$ s to $497.75$ s, IMRPhenomPv2_NRTidalv2 from $0.39$ s to $462.87$ s, and IMRPhenomXP_NRTidalv2 from $0.38$ s to $454.51$ s. These results underscore that without multi-core parallel optimization, traditional methods become notably less efficient compared to the cAE for large-scale waveform generation.
This advantage makes our model particularly suitable for real-time waveform generation and large-scale parameter space sampling, providing a new and feasible approach for efficient GW data analysis.
Figure 9: Comparison of computation time between cAE and traditional waveform approximation. The horizontal axis represents the number of generated samples, while the vertical axis represents the computation time.
Table 2: Waveform generation time (s) for different batch size using the cAE and some waveform approximation methods.
Approximants Batch size 1 10 50 100 500 1000 5000 10000
cAE 0.1188 0.1229 0.1895 0.1954 0.5030 0.8595 3.7589 7.4778
SpinTaylorT1 0.2926 0.8191 3.1774 5.7059 26.0688 51.3171 245.5291 486.8640
IMRPhenomPv2_NRTidal 0.6213 1.5070 3.8230 6.7868 28.2600 53.4714 251.7113 497.7530
IMRPhenomPv2_NRTidalv2 0.3925 0.9189 3.2646 5.6403 23.8314 47.2493 230.3046 462.8673
IMRPhenomXP_NRTidalv2 0.3823 0.9026 3.2442 5.5723 23.6224 46.1181 227.1048 454.5058

### v4

@@ III.2 Architecture and Hyperparameters of the Model
Our study employs the cAEs to model the amplitude and phase of GW waveforms from BNS mergers. The model consists of two independent cAEs, each responsible for learning a low-dimensional representation of either the amplitude $y_{A}$ or phase $y_{\Phi}$ and reconstructing waveforms conditioned on $\Theta$. In our implementation, the dimensionality of the latent representation is set to 300. Each cAE comprises two encoders (Encoder 1 and Encoder 2) and a decoder. Encoder 1 processes the amplitude or phase data, while Encoder 2 encodes the $\Theta$, and the two latent representations are combined in the latent space before being mapped back to a complete waveform by the decoder. The overall model architecture is shown in FIG. 3, where Training and Test sections correspond to the training and inference workflows, while the remaining sections detail the structural components.
Figure 3: Training and testing architecture of our model, along with detailed structures of individual components, the structure of the Transformer and ResNet block are shown in Appendix A (FIG 12).
As shown in FIG. 3, our model consists of two main branches: one for waveform encoding and reconstruction, and one for encoding $\Theta$. The goal is to align latent representations from both branches while ensuring accurate waveform reconstruction.
@@
Latent variables and loss functions: During training, both the waveform and the $\Theta$ are encoded into their respective latent representations, denoted as $z_{A/\Phi}$ and $z_{\Theta}$. To ensure that the decoder can accurately reconstruct the input waveform and that both latent spaces are aligned, we define two loss components: First, the reconstruction loss is computed as the Mean Absolute Error (MAE) between the input waveform $x$ and the reconstructed waveform $\hat{h}$:
@@
The total loss is then defined as a weighted sum of the reconstruction and latent consistency losses, as shown in Eq. 13:
@@ IV Training and Verification
IV.1 Training
Amplitude and phase models are trained separately on an NVIDIA A800 GPU with 80 GB of memory. The full training process, including both models, takes approximately ten days. The training dataset consists of $1\times 10^{6}$ waveform samples, split into 90% for training and 10% for validation. Before training, input parameters and target waveforms are normalized using Min-Max and Standard scaling, respectively. Data is preprocessed into 10 blocks and fed into the model with a batch size of 128. FIG. 4 shows the training and validation loss curves for both amplitude and phase models.
Our model is trained using the Adam optimizer with an initial learning rate of $10^{-4}$. To enhance convergence and prevent stagnation, a learning rate scheduling strategy is employed: if the validation loss does not improve for 7 consecutive epochs, the learning rate is reduced by a factor of 0.7, with a floor value of $10^{-8}$. In addition, early stopping is activated when the validation loss shows no improvement over 15 consecutive epochs, and the model is restored to the state with the best validation performance.
@@ IV.1 Training
IV.2 Verification
To evaluate the precision of cAE-generated waveforms, we compute the mismatch between the model-generated waveforms and the IMRPhenomXP_NRTidalv2 waveforms. The evaluation is based on two metrics: Overlap [55, 56] and Mismatch, which quantify the waveform reconstruction quality. The analysis is conducted in the frequency domain after applying a Fourier transform to the time-domain waveforms, allowing for a more effective comparison of waveform similarities. In GW data analysis, the inner product of two waveforms is typically defined as a noise-weighted integral over the frequency domain, incorporating the power spectral density $S_{n}(f)$:
$\langle h_{1}\mid h_{2}\rangle=4\text{Re}\int_{f_{\min}}^{f_{\max}}\frac{\tilde{h}_{1}(f)\,\tilde{h}_{2}(f)}{S_{n}(f)}\,df,$ (14)
where $\tilde{h}_{1}(f)$ and $\tilde{h}_{2}(f)$ are the Fourier transforms of $h_{1}(t)$ and $h_{2}(t)$, respectively, and $S_{n}(f)$ represents the power spectral density (PSD) of the detector noise. This weighted inner product provides a measure of how well two waveforms match in the presence of detector noise. We set $S_{n}(f)$ to the LIGO O4 sensitivity curve [57].
@@ V Results and Analysis
In this section, we not only compare our proposed conditional autoencoder with traditional waveform approximants, but also with a recent deep learning architecture [58], which employs a residual-stacked Multilayer Perceptron (MLP) and Convolutional Neural Network (CNN). Traditional waveform approximants are included as baseline references in both the mismatch analysis and the generation time comparison. We present the generation precision and generation speed of both the cAE and the residual-stacked MLP-CNN model, using the traditional models as a benchmark. In addition, we further analyze the latent space to examine whether physically meaningful correlations are preserved when varying parameters. Moreover, we perform a tidal phasing consistency test by varying tidal deformabilities $(\Lambda_{1},\Lambda_{2})$ with fixed masses and spins, and compare the cumulative phase evolution against the IMRPhenomXP NRTidalv2 model.
@@ V.1 Mismatch Evaluation
To evaluate the accuracy of waveform reconstruction between neural network models and traditional waveform approximants, we compare our conditional autoencoder model with the Residual-MLP-CNN model, whose architecture is reproduced from [58] and shown in FIG. 13 of Appendix A. Both models are trained using the same dataset and identical preprocessing pipelines, following the same procedure used for our cAE model. The training strategy including optimizer, learning rate scheduler, early stopping, and callbacks—is also kept consistent with that of cAE model. Using the test set introduced in Section II.3, we evaluate waveform reconstruction accuracy using the mismatch. FIG. 5 shows the mismatch distributions of $h_{+}$ and $h_{\times}$ for both models. The average mismatch is calculated as $\text{mismatch}_{\text{avg}}=\frac{1}{N}\sum_{i=1}^{N}\text{mismatch}_{i}$. The average mismatch of the cAE model is $2.13\times 10^{-3}$, while that of the Residual-MLP-CNN model is $5.72\times 10^{-3}$, indicating that the cAE achieves higher reconstruction accuracy.
Figure 5: Mismatch distributions of $h_{+}$ and $h_{\times}$ for both the cAE and Residual-MLP-CNN models over test set.
We analyze the relationship between the number of orbital cycles and the mismatch by computing the cycle count for test set. The time-domain signal $s(t)$ undergoes a Hilbert transform to obtain its analytic representation[59]:
@@
$\text{Cycles}=\frac{\Delta\phi}{2\pi}.$ (21)
FIG. 6 illustrates how the waveform mismatch varies with the number of orbital cycles in the test set. Across a wide range of cycles (200–400), there is no clear trend of mismatch increasing with cycle count. In particular, our cAE exhibits consistently low mismatch values, indicating robust learning performance. The Residual-MLP-CNN model also maintains stable mismatch across the cycle range, though its overall mismatch is higher than that of cAE. These results demonstrate that, given a sufficiently large training dataset ($1\times 10^{6}$ samples), both models are capable of learning accurate waveform representations without degradation at higher cycle counts.
Figure 6: The horizontal axis shows the number of orbital cycles in the test set, and the vertical axis shows the mismatch. Color indicates sample density.
Compared to Residual-MLP-CNN, cAE achieves lower mismatch over nearly the entire cycle range, further highlighting its generalization capacity.
To further evaluate the physical consistency of our cAE model, we conduct a tidal phasing consistency test. In this test, both component masses are fixed at $1.4M_{\odot}$, and the spins are set to zero for both stars, while the tidal deformabilities $(\Lambda_{1},\Lambda_{2})$ are varied. We analyze the cumulative phase evolution $\Phi(t)$ of IMRPhenomXP_NRTidalv2 waveforms under different tidal deformability parameters. Figure 7 shows the cumulative phase $\Phi(t)$ over time for several representative choices of $(\Lambda_{1},\Lambda_{2})$. As the total tidal deformability $\Lambda_{\Sigma}=\Lambda_{1}+\Lambda_{2}$ increases, the resulting phase curves exhibit consistent upward shifts, reflecting the acceleration of phase evolution due to stronger tidal effects.
Figure 7: Cumulative phase $\Phi(t)$ evolution for cAE-generated waveforms under different tidal deformabilities $(\Lambda_{1},\Lambda_{2})$. The inset shows the zoom-in around the merger time.
In FIG. 8, we present the phase difference $\Delta\Phi(t)$ between waveforms generated by our cAE model and the IMRPhenomXP_NRTidalv2. As time increases, the cumulative phase difference gradually grows, reflecting the accumulation of small discrepancies during the inspiral evolution. Nevertheless, the absolute magnitude of $\Delta\Phi(t)$ remains within a relatively small range throughout the entire duration time, indicating that the cAE maintains high accuracy in phase reconstruction. Importantly, when keeping the component masses and spins fixed and varying only the tidal deformabilities $(\Lambda_{1},\Lambda_{2})$, we observe a small upward trend in the mismatch as $\Lambda_{\Sigma}$ increases. This behavior is physically consistent, since stronger tidal effects lead to larger phase shifts in binary neutron star waveforms. However, the mismatch stays at a low level, showing that our cAE can model tidal effects reliably and accurately.
Figure 8: Phase difference $\Delta\Phi(t)$ between the cAE model and IMRPhenomXP_NRTidalv2 over time for different total tidal deformabilities $\Lambda_{\Sigma}$.
To further verify the robustness of the cAE in the strong tidal regime, we examined how the waveform mismatch changes with the total tidal deformability $\Lambda_{\Sigma}=\Lambda_{1}+\Lambda_{2}$, and compared it with the Residual-MLP-CNN model. As shown in FIG. 9, both models exhibit a small increasing trend in mismatch as $\Lambda_{\Sigma}$ grows, which is consistent with the cumulative phase shifts induced by tidal effects. However, the cAE maintains its mismatch at a consistently lower level (below $2\times 10^{-3}$).
Figure 9: Mismatch between cAE and IMRPhenomXP_NRTidalv2 waveforms across different values of total tidal deformability $\Lambda_{\Sigma}=\Lambda_{1}+\Lambda_{2}$. The shaded regions indicate the spread of mismatch values within approximately one standard deviation around the median at each total tidal deformability $\Lambda_{\Sigma}$.
V.2 Latent Space Analysis
To more systematically examine whether the proposed cAE model captures physically meaningful structures within its latent representation, we have added three controlled experiments to , with the corresponding results shown in FIG. 10. Specifically, in each experiment, we vary only a kind of physical parameter while keeping the others fixed to observe the model’s latent vector response: (a) we fix the component masses $m_{1}=m_{2}=1.4,M_{\odot}$ and spins $\chi_{1z}=\chi_{2z}=0$, and gradually increase the tidal deformability $\Lambda_{1}=\Lambda_{2}$; (b) we fix $\Lambda_{1}=\Lambda_{2}=0$ and $\chi_{1z}=\chi_{2z}=0$, and vary the total mass $M_{\mathrm{tot}}$; (c) we fix the masses and tidal parameters, and vary the total spin $\chi_{\mathrm{tot}}$. For each configuration, we input the physical parameters into the cAE and extract the latent vector $z\in\mathbb{R}^{300}$ from the encoder trained on waveform phase.
To compare how each physical variation affects the latent dimensions, we compute the difference $\Delta z=z_{i}-z_{0}$ for each sample relative to the baseline (i.e., the first sample in each set with the smallest parameter value), and sort the latent dimensions by their overall sensitivity. This ensures that dimensions with the most pronounced changes are shown on the left in each figure. Each figure displays 20 samples arranged in a $4\times 5$ grid, where each row contains 5 samples corresponding to different physical configurations. For each sample, we obtain a 300-dimensional latent vector from the encoder (trained on waveform phases). The first sample (top-left corner) serves as the baseline, and the remaining 19 samples are compared against it. Specifically, we compute the absolute difference $\Delta z=|z_{i}-z_{0}|$ between each sample’s latent vector $z_{i}$ and the baseline vector $z_{0}$. Each column represents a latent dimension, and the color intensity reflects the magnitude of $\Delta z$ in that dimension. Brighter colors indicate stronger deviation from the baseline. In addition, a brighter color can be interpreted as a stronger activation of the corresponding latent neuron. A unified color scale is used across all subfigures to ensure consistency in visual interpretation.
The results show interpretable correlations between physical parameters and latent responses. For example, in the tidal variation experiment, only a small subset of latent dimensions exhibits significant and stable increases in activation, suggesting that the cAE has learned to encode tidal effects—which primarily impact late-inspiral phase evolution—into a compact subspace (see Fig. 10(a)). In the mass variation experiment, lower total mass leads to both stronger and more widespread activations, consistent with the fact that lower-mass binaries produce longer inspirals with richer phase features (see Fig. 10(b)). In contrast, for spin variation, the response is more diffuse across many latent dimensions, reflecting the more complex and nonlocal influence of spin, which affects both phase evolution and precession modulation(see Fig. 10(c)).
Figure 10: Latent space analysis of the cAE model under controlled variations of physical parameters. Each subfigure contains 20 samples arranged in a $4\times 5$ grid, where each row contains 5 waveform configurations. Each sample is represented by a horizontal strip of 300 cells, corresponding to the 300-dimensional latent vector output by the encoder. Each column thus represents a specific latent dimension across all samples. The color of each cell indicates the relative activation magnitude $\Delta z$ compared to a baseline (the first sample in each subfigure), with brighter colors representing stronger positive deviations. The baseline row appears uniformly dark by construction. All subfigures share a unified color scale for consistent visual comparison. Experimental setups: (a) Varying tidal deformabilities ($\Lambda_{1}=\Lambda_{2}$) with fixed component masses ($m_{1}=m_{2}=1.4M_{\odot}$) and spins ($\chi_{1z}=\chi_{2z}=0$); (b) Varying total mass with fixed spins ($\chi_{1z}=\chi_{2z}=0$) and tidal deformabilities ($\Lambda_{1}=\Lambda_{2}=0$); (c) Varying total spin $\chi_{\rm tot}$ with fixed component masses and tidal deformabilities.
In Appendix A, FIGs. 16-19 illustrate several examples of reconstructed waveforms generated by our conditional autoencoder model for test cases with different physical parameters.
V.3 Waveform Generation Efficiency
We evaluate waveform generation speed for three categories of models: (i) our conditional autoencoder, (ii) the reproduced Residual-MLP-CNN, (iii) four CPU-based approximants (SpinTaylorT1, IMRPhenomPv2_NRTidal, IMRPhenomPv2_NRTidalv2, and IMRPhenomXP_NRTidalv2). All timings were obtained on an NVIDIA A800 80GB GPU and two Intel Xeon Silver 4214R CPU (24 cores).
Single waveform.
Table 2 shows that our cAE requires 0.1204 s on GPU and 0.2314 s on CPU to generate a single waveform. Compared to traditional CPU-based approximants, the cAE achieves a speed-up of approximately 5–9 times on GPU and remains slightly faster on CPU. When compared to the Residual-MLP-CNN, the two models exhibit nearly identical performance on GPU (0.1204 s vs. 0.1221 s), and similarly on CPU (0.2314 s vs. 0.1675 s), indicating broadly consistent generation times across both architectures. Overall, both neural models significantly outperform traditional approximants in single waveform generation time.
Table 2: Time to generate a single waveform. Entries marked with an asterisk (*) denote AI-based models. GPU-based results are obtained on an NVIDIA A800 80GB, while CPU results are from two Intel Xeon Silver 4214R. Traditional approximants are evaluated on CPU only.
Waveform Time (s)
*cAE (GPU) 0.1204
cAE (CPU) 0.2314
Residual-MLP-CNN (GPU) 0.1221
Residual-MLP-CNN (CPU) 0.1675
SpinTaylorT1 0.9415
IMRPhenomPv2_NRTidal 1.1026
IMRPhenomPv2_NRTidalv2 0.8383
IMRPhenomXP_NRTidalv2 0.6663
Batched generation.
Table 3 presents waveform generation times for batch sizes ranging from 1 to 1000. The conditional autoencoder (cAE) and Residual-MLP-CNN exhibit similar performance across all batch sizes. On GPU, their generation times are nearly identical: at batch size 1, cAE–GPU requires 0.1204 s, compared to 0.1221 s for Residual-MLP–GPU; at batch size 1000, the times are 1.0210 s and 1.0493 s, respectively. On CPU, their runtimes also remain close, for example, at batch size 50, cAE–CPU takes 0.6773 s while Residual-MLP–CPU takes 0.5567 s.
Compared to the traditional approximant IMRPhenomXP_NRTidalv2 running on CPU, both neural models executed on GPU yield speed-ups. For batch size 1, cAE–GPU is approximately 5.54 times faster (0.1204 s vs. 0.6663 s), and for batch size 1000, the speed-up increases to 9.92 times (1.0210 s vs. 10.1279 s). Residual-MLP–GPU shows a similar trend, reaching 9.65 times faster at batch size 1000.
Table 3: Waveform generation time for different batch sizes. Residual-MLP-CNN is abbreviated as ResMC. ResMC–GPU and ResMC–CPU refer to ResMC evaluated on an A800 GPU and on two Xeon Silver 4214R CPUs, respectively. Similarly, cAE–GPU and cAE–CPU denote our conditional autoencoder model evaluated on GPU and CPU. IMR–CPU refers to IMRPhenomXP_NRTidalv2 evaluated on the same CPU system. The speed-up ratios are defined as SU(IMR/cAE-G) = t(IMR-CPU) / t(cAE-GPU) and SU(IMR/ResMC-G) = t(IMR-CPU) / t(ResMC-GPU), where t denotes the time.
Batch size cAE–GPU cAE–CPU ResMC–GPU ResMC–CPU IMR–CPU SU(IMR/cAE–G) SU(IMR/ResMC–G)
1 0.1204 s 0.2314 s 0.1221 s 0.1675 s 0.6663 s 5.54 5.46
10 0.1500 s 0.3235 s 0.1318 s 0.2797 s 0.7113 s 4.74 5.40
50 0.1663 s 0.6773 s 0.1726 s 0.5567 s 1.0646 s 6.40 6.17
100 0.2501 s 1.2440 s 0.2291 s 1.1831 s 1.1377 s 4.55 4.97
500 0.5902 s 5.8146 s 0.5927 s 5.5440 s 5.2263 s 8.86 8.82
1000 1.0210 s 11.7420 s 1.0493 s 11.1276 s 10.1279 s 9.92 9.65
Bayesian inference efficiency.
To further demonstrate the practical utility of our cAE model in downstream applications, we performed a simplified Bayesian inference experiment on the binary neutron star masses $(m_{1},m_{2})$. The injected (true) source parameters were set to $m_{1}^{\rm true}=1.45\,M_{\odot}$, $m_{2}^{\rm true}=1.25\,M_{\odot}$, with dimensionless spins $\chi_{1z}=\chi_{2z}=0.1$ and tidal deformabilities $\Lambda_{1}=\Lambda_{2}=200$. The detector noise was modeled using the aLIGO O4 sensitivity curve, and posterior sampling was performed with the emcee package[60]. We employed 4 chains of 5000 steps each, discarding the first 1000 steps as burn-in. Since the purpose of this experiment is primarily to compare the efficiency of waveform templates, rather than to perform full high-dimensional parameter estimation, we fixed the spin and tidal parameters and inferred only $(m_{1},m_{2})$. This simplification is sufficient to reflect the computational cost and efficiency differences between traditional approximants and the deep-learning cAE templates.
The inference results are shown in FIG. 11. For the cAE template, the estimated parameters are $m_{1}=1.469$ with a standard deviation of $0.075$ ($m_{1}=1.469\pm 0.075\,M_{\odot}$) and $m_{2}=1.245$ with a standard deviation of $0.060$ ($m_{2}=1.245\pm 0.060\,M_{\odot}$), with 68% credible intervals of $(1.387,1.550)$ and $(1.178,1.312)$, respectively. For the traditional approximant, the estimates are $m_{1}=1.530$ with a standard deviation of $0.078$ ($m_{1}=1.530\pm 0.078\,M_{\odot}$) and $m_{2}=1.191$ with a standard deviation of $0.059$ ($m_{2}=1.191\pm 0.059\,M_{\odot}$), with 68% credible intervals of $(1.427,1.607)$ and $(1.134,1.353)$, respectively.
The results suggest that the cAE model delivers parameter estimates with accuracy on par with the IMRPhenomXP_NRTidalv2 approximant. In terms of runtime, the cAE model provides a meaningful improvement in computational efficiency. Under the same sampling setup, the traditional approximant required approximately 312 seconds using 24 CPU cores, while the cAE completed the inference in only 61 seconds on a single NVIDIA A800 GPU—resulting in a speedup of approximately 5.1 times. These findings provide initial evidence that the cAE model not only maintains high waveform reconstruction fidelity but also offers significant computational benefits in practical inference workflows. This makes it particularly well-suited for time-sensitive applications such as rapid alerts or large-scale population analyses. We hope this experiment offers a useful reference point for future extensions to more complex, higher-dimensional inference scenarios.
Figure 11: Posterior distributions of binary neutron star component masses $(m_{1},m_{2})$ obtained from Bayesian inference.